A rigid container holds 1 kg of air initially at 4.8 bar and 150 degree Celsius the air is heated until its temperature is 200 degree Celsius determine the work done
1
Expert's answer
2021-09-06T00:49:21-0400

Mass of air = 1 kg

Initial temperature T1 = 150 + 273 = 423 K

Initial pressure P1 = 4.8 bar

Final temperature T2 = 200 + 273 = 473 K

The container is rigid (volume is constant)

From the ideal gas equation


"P2=P1*T2\/T1"

=4.8bar(473/423)

=5.37 bars

For isochoric process

Entropy change "S = Cv ln(\frac{T2}{T1})"

​= 0.08 kJ/kgK


Average temperature T = "\frac{(473 + 423)}{2} = 448 K"

Work done= "T \times S"

"= 448 K x 0.08 kJ\/kgK\n\n= 35.9 kJ\/kg"